using ages 12, 40, and 38 write an equation that models how old in years each of you will be, when your ages add up to 150 years old
(y + a) + (y + a) + (y + a) = 150 yrs.
(12 + a) + (40 + a) + (38 + a) = 150 yrs.
90 + a3 = 150 yrs.
This is what I came up with is it right?

Answer by Theo(13342)

current ages are 12, 40, and 38
you want to know in how many years the sum of these ages will be 150.
let y equal the number of years from now.
your equation becomes:
12 + y + 40 + y + 38 + y = 150
this becomes:
90 + 3y = 150
subtract 90 from both sides of this equation to get:
3y = 60
divide both sides of this equation by 3 to get:
y = 20
in 20 years, the sum of your ages will be 150.
12 + 20 = 32
40 + 20 = 60
38 + 20 = 58
32 + 60 + 58 = 150
your equation didn't look quite right but you got the right answer so you must have been thinking correctly even though someone who looked at your equation might not have thought so in the beginning.
your problem in your equation is you were using y to represent 3 different ages at one time.
normally that would be shown as 3 different variables.
assume the following:
let y1 = the age of the first person
let y2 = the age of the second person
let y3 = the age of the third person
let a = the number of years
then your equation would have become:
(y1 + a) + (y2 + a) + (y3 + a) = 150
this would then have been simplified to:
y1 + y2 + y3 + 3a = 150
you could then substitute 12 for y1 and 40 for y2 and 38 for y3 to get:
12 + 40 + 38 + 3a = 150 which would have led to:
90 + 3a = 150
you would then have solved as before using standard equation solving procedures.